OpenStudy (anonymous):

Solve: 1) 3^(8-2x)=81 2) ln 4x + ln 3x = ln 13

OpenStudy (anonymous):

1) 3 to what power is 81? Then 8-2x must be equal to that number. 2) use the properties of logs to rewrite the LHS (Left Hand Side). then eponentiate each side and solve.

OpenStudy (anonymous):

1) =4 2)= i have no clue

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

1) 4 = 8-2x. solve for x. 2) ln x + ln y = ln xy use this property. then ln xy = ln z => xy = z. solve.

OpenStudy (anonymous):

so 1) is 2? and 2) I'm working on

OpenStudy (anonymous):

where did z come from? not sure what to plug in for that

OpenStudy (hitaro9):

z is arbitrary, he's just using it as an example.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh okay. so to be honest i am utterly confused with this log stuff

OpenStudy (hitaro9):

Okay. You'll get a better understanding of it as you study more. For right now you just need to know specific processes. You know that ln(xy) = ln(x) +ln(y) right?

OpenStudy (anonymous):

yes, if i use EXACT numbers. but what do i do with the numbers after finding the ln of them?

OpenStudy (anonymous):

ln 3x + ln 4x = ln (3x)(4x) = ln (12x^2)

OpenStudy (hitaro9):

Well with this problem you don't need to find the ln of them.

OpenStudy (hitaro9):

As pgpilot has just shown, ln(3x)+ln(4x) =ln(12x^2)

OpenStudy (hitaro9):

You therefore know that ln(12x^2) = ln(13)

OpenStudy (anonymous):

so 12x^2 = 13 solve for x.

OpenStudy (hitaro9):

^ Right

OpenStudy (hitaro9):

If you see two things in the same "situation" in an equation, you can remove the rest of the stuff So like, 3^x = 3^4 you can assume that x=4 without having to deal with logs or anything complicated like that

OpenStudy (hitaro9):

Similarly, if you see that ln(x) = ln(2) you can just assume that x is equal to 2

OpenStudy (hitaro9):

Apply the same thing to your equation ln(12x^2) = ln(13) can be changed to 12x^2 = 13 And from there it's basic algebra.

OpenStudy (anonymous):

bnut remember that whatever you find x is, when it goes back in to the origianl equation, the argument of log or ln must be positive. in ln x x is the argument. so x > 0.

OpenStudy (anonymous):

its basic but it comes out as a huge decimal. is positive and−1.040832999733what they want?

OpenStudy (hitaro9):

You don't need to worry about that

OpenStudy (anonymous):

\[x = \pm \sqrt{\frac{13}{12}}\] but only the positive value is valid

OpenStudy (anonymous):

thank you guys:)

OpenStudy (hitaro9):

Yeah. That ^ Your teacher doesn't need to see 1,000,000 digits you got out of your calculator.

OpenStudy (anonymous):

up for anymore questions?

OpenStudy (hitaro9):

If you just apply the same principles, I'm sure it'll work for the rest of your homework.

OpenStudy (anonymous):

yeah, put it in a new post. okay?

OpenStudy (hitaro9):

But go ahead

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

no worries!